Convert the following units of Force:

1 Avogram•barn/(hubble•millennium2) to Exawatt•millennium/zeptometer
Answer #1

avogram = 1gm/6.023*10^23 = 0.17*10^-23 gm

1 barn = 10^-24 cm^2

1 milinium = 3.15*10^10 sec

1 hubble = 1.3*10^26 m

1 Avogram•barn/(hubble•millennium2) = 0.013*10^-100 N

1 exa watt = 10^18 watt

1 zepto meter = 10^-12 m

1 millennium = 3.15*10^10 sec

Exawatt•millennium/zeptometer = 3.15*10^40 N